A busy chipmunk runs back and forth along a straight line of acorns that has been set out between its burrow and a nearby tree. At some instant, it moves with a velocity of -1.31 m/s. Then, 2.03 s later, it moves with a velocity of 1.81 m/s. What is the chipmunks average acceleration during the 2.03 s time interval? average acceleration: m/s2
Answer #1

Acceleration = v-u/t u = -1.31m/s v= 1.81 m/s t= 2.03

a= 1.81-(-1.31)/2.03= 3.12/2.03=1.53 m/s^2

average acceleration = 1.53 m/s^2